The formula s=-16t^2 + vt + s gives you the height in feet above the earth at time t seconds for an object projected into the air with an initial velocity of v feet per second (ft/sec) from an initial height of s feet. Two balls are thrown into the sir simultaneously, one from the ground at 40 ft/sec and one from a height of 20 feet at 30 ft/sec. For what values of t will their heights above the ground differ by 10 feet or less?
Answer by stanbon(75887)

The formula s=-16t^2 + vt + s gives you the height in feet above the earth at time t seconds for an object projected into the air with an initial velocity of v feet per second (ft/sec) from an initial height of s feet.
-----------------------------------------------------------------------
Two balls are thrown into the sir simultaneously, one from the ground at 40 ft/sec and one from a height of 20 feet at 30 ft/sec.
1st ball: s(t) = -16t^2+40t+0
2nd ball: s(t) = -16t^2+30t+20
---------------------------------------
For what values of t will their heights above the ground differ by 10 feet or less?
-------------------------------------
Equation:
height - height <= 10 feet
(-16t^2+30t+20) - (-16t^2+40t) <= 10
---
-10t+20 <= 10
-10t <= -10
t >= 1 second (the heights will differ by at least 10 ft after the 1st second
and until one of the balls reaches the ground.
===================================================
Cheers,
Stan H.